# Rotor-Failure-Aware Quadrotors Flight in Unknown Environments

### v1

@@ I Introduction
The second challenge lies in ensuring reliable perception and collision avoidance in unknown environments, while accommodating altered dynamic feasibility following rotor failures. To maintain thrust balance, the thrust of the rotor opposite to the failed one is restricted to a lower amplitude [12]. This constraint significantly alters the quadrotor’s thrust-to-gravity ratio and yaw torque, changing the safe flight envelope and dynamic feasibility [13]. Consequently, trajectory planning must account for the change in dynamic feasibility before and after rotor failures. Additionally, the imbalance in yaw torque induces high-speed rotation and vibration, severely degrading the perception performance of autonomous navigation systems [14]. As a result, many related studies rely on global positioning systems (GPS) or motion capture systems (MCS) for post-failure trajectory tracking [15, 12], limiting their applicability in unknown and complex environments. In scenarios lacking global information and featuring unknown obstacles, global trajectory tracking without real-time perception cannot ensure flight safety. Addressing the second challenge requires a reliable perception and trajectory planning system that accounts for high-speed rotation and altered dynamic feasibility.
@@ III-A Notations
Throughout this paper, the set of real numbers is denoted as $\mathbb{R}$, while $\text{diag}[d_{1},d_{2},d_{3}]$ refers to a diagonal matrix with $d_{1}$, $d_{2}$, and $d_{3}$ on its diagonal. For a matrix $\bm{A}\in\mathbb{R}^{m\times n}$ or a vector $\bm{x}\in\mathbb{R}^{n}$, the transpose is denoted by $\bm{A}^{\top}$ or $\bm{x}^{\top}$, respectively. The Euclidean norm of $\bm{x}$ is written as $\|\bm{x}\|$, defined by $\|\bm{x}\|=\sqrt{\bm{x}^{\top}\bm{x}}$. The coordinate systems of the quadrotor consist of two right-handed frames: the inertial frame, $F_{w}:\{\bm{x}_{w},\bm{y}_{w},\bm{z}_{w}\}$, with $\bm{z}_{w}$ pointing upward, opposite to gravity; and the body frame, $F_{B}:\{\bm{x}_{B},\bm{y}_{B},\bm{z}_{B}\}$, where $\bm{x}_{B}$ points forward and $\bm{z}_{B}$ aligns with the thrust direction (see Fig. 2). The matrix and vector with the subscript $B$ are represented in the body frame, while the matrix and vector without a subscript are represented in the inertial frame. The transformation from $F_{w}$ to $F_{B}$ is defined by the rotation matrix $\bm{R}$, which is parameterized by the quaternion $\bm{q}=[q_{w},q_{x},q_{y},q_{z}]^{T}$.
@@
$m\in\mathbb{R},\bm{g}\in\mathbb{R}^{3}$ Mass of the quadrotor and gravity vector.
$\bm{t}\in\mathbb{R}^{3}$ Thrust vector of the quadrotor.
@@
$\kappa_{t}\in\mathbb{R}$ The yaw torque coefficient.
@@ III-B Quadrotor Dynamics
where $\bm{\eta}$ represents the position of the quadrotor’s center of gravity (CoG). $\bm{v}$, $T$, $\bm{D}$, $m$, and $\bm{g}$ denote the velocity vector, total thrust, rotor-drag coefficient vector, the mass, and the gravity vector. In the rotational dynamics, $\otimes$ denotes quaternion multiplication, $\bm{w}_{B}=[p,q,r]^{T}$ represents the angular velocity of frame $\mathcal{F}_{B}$ relative to $\mathcal{F}_{w}$, and $\bm{\alpha}_{B}$ is the corresponding angular acceleration. The term $\bm{I}_{v}$ denotes the quadrotor’s inertia matrix, while $\bm{\tau}=[\tau_{x},\tau_{y},\tau_{z}]^{T}$ is the total torque generated by the rotors. Additionally, $\bm{A}(r)=[0,\ 0,\ -k_{d,\psi}r]^{T}$ represents the torque induced by rotation motion, and $k_{d,\psi}$ is the aerodynamic yaw damping coefficient. Details of the variables and parameters are listed in the Table I. It is noted that the above dynamic model does not capture certain unmodeled effects such as aerodynamic uncertainties, wind disturbance, and sensor noise. While these factors may affect state estimation and control accuracy, the proposed framework mitigates them through the integration of LiDAR-inertial odometry and NMPC.
@@ IV-A Software Framework
To enable autonomous quadrotors flight under rotor failures, the software framework illustrated in Fig. 3, comprising perception, planning, and control modules, is integrated into the onboard computer. The onboard computation is performed on an NVIDIA Jetson Orin NX, a powerful embedded platform with an 8-core CPU running Ubuntu 20.04. This hardware provides sufficient computational capability to support real-time NMPC and perception modules. Communication between the onboard computer and the flight controller is facilitated via MAVROS [29]. The quadrotor’s software system primarily consists of the following subsystems.
@@ IV-B Quadrotor Platform Design
All experiments are performed on a custom-designed and assembled quadrotor platform with a wheelbase of 250 $mm$. The hardware specifications of this platform are presented in Fig. 4. The platform weighs 1150 $g$ and measures 31.6 × 31.6 × 18.2 $cm$. It is powered by a 2300 $mAh$ lithium-polymer battery and equipped with four 2550 $KV$ brushless motors and 5-inch three-blade propellers. When a rotor failure occurs, the imbalance in yaw torque causes the quadrotor to enter a high-speed spinning motion accompanied by intense vibrations. Under these conditions, the efficiency of state estimation and mapping significantly deteriorates. First, the inertial measurement unit (IMU) is subjected to substantial centrifugal acceleration, resulting in increased noise and errors in the measurement of acceleration and angular velocity [14]. Second, the odometry accuracy degrades due to motion blur and scan distortions caused by high-speed motion [28]. The FAST-LIO2 algorithm [30], an efficient and robust LiDAR-inertial odometry framework, is utilized. Besides, the mechanical design and sensor selection contribute to meeting the demands of localization and mapping during the high-speed self-rotation of the post-failure quadrotors.
@@
where $\tau_{z}$ denotes the yaw torque acting on the quadrotor. In the event of rotor failures, the imbalance in $\tau_{z}$ leads to an increase in the spin speed $r$ until the air resistance generated by rotation, $k_{d,\psi}r$, balances $\tau_{z}$. To mitigate excessive spin speed, a viable strategy is to adjust the torque coefficient $k_{t}$ and the air resistance coefficient $k_{d,\psi}$. Therefore, we improve the quadrotor platform with four anti-torque plates mounted on the landing gears. These plates increase air resistance and compensate for the motor’s counter-torque. The anti-torque plates have a width of 35 $mm$ and a length of 77 $mm$. The plates have a large projected area in the vertical plane coinciding with the arm, which generates large air resistance during quadrotor spinning, thereby increasing the air resistance coefficient $k_{d,\psi}$. Moreover, the airfoil design allows for creating torque in the opposite direction of the propeller’s counter-torque, thereby reducing the torque coefficient $k_{t}$. The torque coefficient $k_{t}$ is obtained from a power test using a professional test bench, provided by Shenzhen Deep Exploration Technology Co., Ltd (http://www.szde-tech.com/). The yaw damping coefficient $k_{d,\phi}$ is identified from rotor-failure flight experiments. These procedures ensure that both parameters reflect the actual quadrotor dynamics. The Livox MID-360 LiDAR and an IMU are selected as the primary sensors (see Fig. 4). The perception under high-speed rotation is alleviated through the 360° horizontal field-of-view (FoV) of the Livox MID-360 LiDAR for omni-directional environmental perception. Simultaneously, the IMU operates at a frequency of 333 Hz, providing estimates of the quadrotor’s acceleration and angular velocity. Data from both the LiDAR and IMU are processed by an onboard extended Kalman filter (EKF).
@@ V-A NMPC Design
By defining the quadrotor state as $\bm{x}=[\bm{\eta}^{T},\bm{q}^{T},\dot{\bm{\eta}}^{T},\bm{w}^{T},\bm{t}^{T}]^{T}$ and the control input as $\bm{u}=[u_{1},u_{2},u_{3},u_{4}]^{T}$, the quadrotor dynamics can be expressed in discrete-time form as $\bm{x}_{k+1}=f(\bm{x}_{k},\bm{u}_{k})$. The NMPC framework seeks to minimize a cost function defined by the errors between the reference and predicted states over a finite-time horizon $t_{N}\in[t,t+h]$. The step size is determined by $dt=h/N$, where $N$ represents the number of time steps and $h$ is the horizon length. The constrained optimization problem can be formulated as:
@@
$\bm{y}_{i}=\left[\begin{matrix}\bm{\eta}-\bm{\eta}_{ref}\\ \tilde{\bm{q}}_{e}\\ \bm{v}-\bm{v}_{ref}\\ \bm{w}-\bm{w}_{ref}\\ \bm{t}-\bm{t}_{ref}\end{matrix}\right],$ (10)
@@
The reference values in the cost functions are derived either from hover references, applied uniformly across all time points in the prediction horizon, or from sampled points along reference trajectories. Under rotor-failure conditions, the weights associated with the yaw channel are set to zero, while the remaining weights are preserved.
Remark 1.
Although the rotor-failure-aware control scheme is only validated on the quadrotor, the proposed method can be extended to other configurations of multirotors. The primary differences between various multirotors lie in the mixing control matrix and the number of motors. By modifying the model information in the FDD and FTC, the proposed fault-tolerant system can be adapted for other multirotors.
@@ V-B Composite FDD for Different Flight Stages
For motor failure diagnosis, we utilize the real-time measured RPM $\hat{w}_{i}$ and compare it against the reference RPM ${w}_{i}$. Motor failure is assessed using the following index:
@@
When $M_{i}\leq\gamma_{M}$ is satisfied where $0<\gamma_{M}\leq 1$ serves as the threshold for determining motor failure, the control mode switches to FTC.
@@
When $P_{i}<\gamma_{P}$ where $0<\gamma_{P}\leq 1$ serves as the trigger threshold, the $i$th propeller is classified as failed and isolated. The failure information of the $i$th propeller is subsequently transmitted to trigger controller reconfiguration and trajectory planning.
@@
Beyond complete rotor failures, the FDD framework can also detect the partial thrust reductions (e.g., due to wear, overheating) by combining a motor-side cue and a propeller-side cue. Intermittent degradations are detectable if they persist within a short observation window or appear with sufficient frequency. Certain mechanical anomalies such as shaft misalignment typically manifest as speed–thrust inconsistencies and are therefore also within the detectable scope of the proposed FDD. Extremely subtle or very short-lived anomalies may remain undetected, suggesting future extensions that incorporate vibration or thermal cues.
Remark 2.
In this paper, FDD for propeller failure during trajectory tracking is a model-based FDD method with dual-loop observers, while the other two FDD methods are model-free approaches that directly utilize the state response information. Integrating model-based and model-free FDD methods into a composite FDD system is essential to ensure timely failure detection despite various types of rotor failures among the takeoff and trajectory tracking stages. This integration provides sufficient reconfiguration time before the post-failure quadrotor enters irreversible status, as compared to prior studies [8, 12, 19].
@@ V-C Rotor-Failure-Aware Trajectory Planning
For a polynomial trajectory with $M=5$ segments in $m$ dimensions and a degree N=2s1N=2s-1 (where $s=3$ is the integrator chain order), a linear-complexity mapping is defined as:
@@
where $\bm{c}=(\bm{c}_{1}^{T},\ldots,\bm{c}_{M}^{T})^{T}\in\mathbb{R}^{2Ms\times m}$ represents the polynomial coefficient matrix. The mapping $M(\bm{q}^{p},\bm{T}^{p})$ smoothly relates the intermediate waypoints $\bm{q}^{p}=(q_{1},\ldots,q_{M-1})^{T}\in\mathbb{R}^{m\times(M-1)}$ and time allocation $\bm{T}^{p}=(T_{1},\ldots,T_{M})^{T}\in\mathbb{R}_{>0}^{M}$ to $\bm{c}$. Using $M(\bm{q}^{p},\bm{T}^{p})$, the optimization variables are transformed into $\{\bm{q}^{p},\bm{T}^{p}\}$. The overall trajectory $p(t)$ and its individual segment $p_{g}(t)$ are expressed as:
$p(t)=p_{g}(t-t_{g-1}),\quad\forall t\in[t_{g-1},t_{g}],$ (20)
$p_{g}(t)=c_{i}^{T}\beta(t),\quad\forall t\in[0,T_{g}],$ (21)
where $\bm{\beta}(t)=[1,t,\ldots,t^{N}]^{T}$ denotes the natural polynomial basis. The trajectory generation for the quadrotor is formulated as an unconstrained nonlinear optimization problem:
@@
We impose limits on linear velocity, acceleration, and jerk to ensure dynamic feasibility during trajectory generation. The corresponding cost function is defined as:
@@
$J_{d,v}=\sum_{g=1}^{k}\max\left\{{\dot{p}^{2}_{g}(t)}-v_{{max}}^{2},0\right\},$ (26)
$J_{d,a}=\sum_{g=1}^{k}\max\left\{{\ddot{p}^{2}_{g}(t)}-a_{{max}}^{2},0\right\},$ (27)
@@
$J_{d,j}=\sum_{g=1}^{k}\max\left\{{\dddot{p}^{2}_{g}(t)}-j_{{max}}^{2},0\right\},$ (29)
@@
To avoid flight accident, the thrust of the rotor opposite to the faulty one is restricted to a small amplitude, reducing the available thrust and torque. If the quadrotor attempts to follow a predefined trajectory without adjusting acceleration, it may exceed its operational limits, jeopardizing flight safety. Therefore, a detailed analysis is performed to balance the post-failure acceleration $a_{{max},f}$ with the remaining thrust.
@@
$a_{{max},f}=\gamma_{a,f}\left|\frac{\sqrt{T_{f,{max}}^{2}-(mg)^{2}}-f_{{max}}}{m}\right|,$ (33)
where $0<\gamma_{a,f}\leq 1$ is a safety factor to enhance flight reliability. To ensure safe operation after rotor failures, we explicitly define dynamic feasibility restrictions, including velocity, acceleration, and jerk constraints, and reduced acceleration capacity (see Eq. 33). These restrictions are incorporated into the optimization formulation to guarantee that generated trajectories remain within the post-failure actuation limits. Although we do not explicitly model efficiency losses or battery degradation, our trajectory optimization already includes total time minimization, which serves as a proxy for energy consumption. This reflects the fact that shorter flight times typically correspond to reduced power usage.
@@
$(\mathbf{x}^{p}-\mathbf{s}^{p})^{\top}\mathbf{v}^{p}=0,\quad\mathbf{x}^{p}\in\mathbb{R}^{3},$ (34)
where $\mathbf{s}\in\mathbb{R}^{3}$ denotes a point on the plane, and $\mathbf{v}^{p}\in\mathbb{R}^{3}$ is the normal vector pointing toward the free space. To quantify the relationship between a given point $\mathbf{p}^{p}\in\mathbb{R}^{3}$ and the obstacle, the distance $d_{o}$ to the plane is computed as:
$d_{o}=(\mathbf{p}^{p}-\mathbf{s}^{p})^{\top}\mathbf{v}^{p}.$ (35)
@@
$J_{c}=\sum_{i=0}^{\kappa}\max\left\{\left(C_{o}-d_{o}(p(t_{g}))\right),0\right\}^{3},$ (36)
where $C_{o}$ denotes safe distance and $d_{o}$ depends on the position.
@@
Figure 6: Simulation visualization with the proposed method in Test #1, where the propeller-ejection failure occurs during quadrotor hovering. The green and red paths indicate the reference lemniscate trajectory and the real flight trajectory, respectively.

### v2

@@ I Introduction
The second challenge lies in ensuring reliable perception and collision avoidance in unknown environments, while accommodating altered dynamic feasibility following rotor failures. To maintain thrust balance, the thrust of the rotor opposite to the failed one is restricted to a lower amplitude [12]. This constraint significantly alters the quadrotor’s thrust-to-weight ratio and yaw torque, changing the safe flight envelope and dynamic feasibility [13]. Consequently, trajectory planning must account for the change in dynamic feasibility before and after rotor failures. Additionally, the imbalance in yaw torque induces high-speed rotation and vibration, severely degrading the perception performance of autonomous navigation systems [14]. As a result, many related studies rely on global positioning systems (GPS) or motion capture systems (MCS) for post-failure trajectory tracking [15, 12], limiting their applicability in unknown and complex environments. In scenarios lacking global information and featuring unknown obstacles, global trajectory tracking without real-time perception cannot ensure flight safety. Addressing the second challenge requires a reliable perception and trajectory planning system that accounts for high-speed rotation and altered dynamic feasibility.
@@ III-A Notations
Throughout this paper, the set of real numbers is denoted as $\mathbb{R}$, while $\text{diag}[d_{1},d_{2},d_{3}]$ refers to a diagonal matrix with $d_{1}$, $d_{2}$, and $d_{3}$ on its diagonal. For a matrix $\bm{A}\in\mathbb{R}^{m\times n}$ or a vector $\bm{x}\in\mathbb{R}^{n}$, the transpose is denoted by $\bm{A}^{T}$ or $\bm{x}^{T}$, respectively. The Euclidean norm of $\bm{x}$ is written as $\|\bm{x}\|$, defined by $\|\bm{x}\|=\sqrt{\bm{x}^{T}\bm{x}}$. The coordinate systems of the quadrotor consist of two right-handed frames: the inertial frame, $F_{w}:\{\bm{x}_{w},\bm{y}_{w},\bm{z}_{w}\}$, with $\bm{z}_{w}$ pointing upward, opposite to gravity; and the body frame, $F_{B}:\{\bm{x}_{B},\bm{y}_{B},\bm{z}_{B}\}$, where $\bm{x}_{B}$ points forward and $\bm{z}_{B}$ aligns with the thrust direction (see Fig. 2). The matrix and vector with the subscript $B$ are represented in the body frame, while the matrix and vector without a subscript are represented in the inertial frame. The transformation from $F_{w}$ to $F_{B}$ is defined by the rotation matrix $\bm{R}$, which is parameterized by the quaternion $\bm{q}=[q_{w},q_{x},q_{y},q_{z}]^{T}$.
@@
$m_{0}\in\mathbb{R},\bm{g}\in\mathbb{R}^{3}$ Mass of the quadrotor and gravity vector.
$\bm{t}\in\mathbb{R}^{4}$ Thrust vector of the quadrotor.
@@
$k_{t}\in\mathbb{R}$ The yaw torque coefficient.
@@ III-B Quadrotor Dynamics
where $\bm{\eta}$ represents the position of the quadrotor’s center of gravity (CoG). $\bm{v}$, $T$, $\bm{D}$, $m_{0}$, and $\bm{g}=[0,0,-g_{0}]^{T}$ denote the velocity vector, total thrust, rotor-drag coefficient, the mass, and the gravity vector. In the rotational dynamics, $\otimes$ denotes quaternion multiplication, $\bm{w}_{B}=[p,q,r]^{T}$ represents the angular velocity of frame $\mathcal{F}_{B}$ relative to $\mathcal{F}_{w}$, and $\bm{\alpha}_{B}$ is the corresponding angular acceleration. The term $\bm{I}_{v}$ denotes the quadrotor’s inertia matrix, while $\bm{\tau}=[\tau_{x},\tau_{y},\tau_{z}]^{T}$ is the total torque generated by the rotors. Additionally, $\bm{A}(r)=[0,\ 0,\ k_{d,\psi}r]^{T}$ represents the torque induced by rotation motion, and $k_{d,\psi}$ is the aerodynamic yaw damping coefficient. Details of the variables and parameters are listed in the Table I. It is noted that the above dynamic model does not capture certain unmodeled effects such as aerodynamic uncertainties, wind disturbance, and sensor noise. While these factors may affect state estimation and control accuracy, the proposed framework mitigates them through the integration of LiDAR-inertial odometry and NMPC.
@@ IV-A Software Framework
To enable autonomous quadrotors flight under rotor failures, the software framework illustrated in Fig. 3, comprising perception, planning, and control modules, is integrated into the onboard computer. The onboard computation is carried out on an NVIDIA Jetson Orin NX, a powerful embedded and edge computing platform equipped with an 8-core CPU, a 1024-core GPU, and 16 GB of RAM. This hardware runs Ubuntu and ROS, and provides sufficient computational power for real-time NMPC and perception modules. Communication between the onboard computer and the flight controller is facilitated via MAVROS [29]. The quadrotor’s software system primarily consists of the following subsystems.
@@ IV-B Quadrotor Platform Design
All experiments are performed on a custom-designed and assembled quadrotor platform with a wheelbase of 250 $mm$. The hardware specifications of this platform are presented in Fig. 4. The platform weighs 1150 $g$ and measures 31.6 × 31.6 × 18.2 $cm$. It is powered by a 2300 $mAh$ lithium-polymer battery and equipped with four 2550 $KV$ brushless motors and 5-inch three-blade propellers. When a rotor failure occurs, the imbalance in yaw torque causes the quadrotor to enter a high-speed spinning motion accompanied by intense vibrations. Under these conditions, the efficiency of state estimation and mapping significantly deteriorates. First, the IMU is subjected to substantial centrifugal acceleration, resulting in increased noise and errors in the measurement of acceleration and angular velocity [14]. Second, the odometry accuracy degrades due to motion blur and scan distortions caused by high-speed motion [28]. The FAST-LIO2 algorithm [30], an efficient and robust LiDAR-inertial odometry framework, is utilized. Besides, the mechanical design and sensor selection contribute to meeting the demands of localization and mapping during the high-speed self-rotation of the post-failure quadrotors.
@@
where $\tau_{z}$ denotes the yaw torque acting on the quadrotor. In the event of rotor failures, the imbalance in $\tau_{z}$ leads to an increase in the spin speed $r$ until the air resistance generated by rotation, $k_{d,\psi}r$, balances $\tau_{z}$. To mitigate excessive spin speed, a viable strategy is to adjust the torque coefficient $k_{t}$ and the air resistance coefficient $k_{d,\psi}$. Therefore, we improve the quadrotor platform with four anti-torque plates mounted on the landing gears. These plates increase air resistance and compensate for the motor’s counter-torque. The anti-torque plates have a width of 35 $mm$ and a length of 77 $mm$. The plates have a large projected area in the vertical plane coinciding with the arm, which generates large air resistance during quadrotor spinning, thereby increasing the air resistance coefficient $k_{d,\psi}$. Moreover, the airfoil design allows for creating torque in the opposite direction of the propeller’s counter-torque, thereby reducing the torque coefficient $k_{t}$. The torque coefficient $k_{t}$ is obtained from a power test using a professional test bench, provided by Shenzhen Deep Exploration Technology Co., Ltd (http://www.szde-tech.com/). The yaw damping coefficient is identified from rotor-failure flight experiments. These procedures ensure that both parameters reflect the actual quadrotor dynamics. The Livox MID-360 LiDAR and an IMU are selected as the primary sensors (see Fig. 4). The perception under high-speed rotation is alleviated through the 360° horizontal field-of-view (FoV) of the Livox MID-360 LiDAR for omni-directional environmental perception. Simultaneously, the IMU operates at a frequency of 333 Hz, providing estimates of the quadrotor’s acceleration and angular velocity. Data from both the LiDAR and IMU are processed by an onboard extended Kalman filter (EKF).
@@ V-A NMPC Design
By defining the quadrotor state as $\bm{x}=[\bm{\eta}^{T},\bm{q}^{T},\dot{\bm{\eta}}^{T},\bm{w}^{T},\bm{t}^{T}]^{T}$ and the control input as $\bm{u}=[u_{0},u_{1},u_{2},u_{3}]^{T}$, the quadrotor dynamics can be expressed in discrete-time form as $\bm{x}_{k+1}=f(\bm{x}_{k},\bm{u}_{k})$. The NMPC framework seeks to minimize a cost function defined by the errors between the reference and predicted states over a finite-time horizon $t_{N}\in[t,t+h]$. The step size is determined by $dt=h/N$, where $N$ represents the number of time steps and $h$ is the horizon length. The constrained optimization problem can be formulated as:
@@
$\bm{y}_{i}=\left[\begin{matrix}\bm{\eta}-\bm{\eta}_{ref},\tilde{\bm{q}}_{e},\bm{v}-\bm{v}_{ref},\bm{w}-\bm{w}_{ref},\bm{t}-\bm{t}_{ref}\end{matrix}\right]^{T},$ (10)
@@ V-B Composite FDD for Different Flight Stages
For motor failure diagnosis, we employ the real-time motor speed $\hat{\omega}_{i}$ reported by the ESC via DShot telemetry. The estimated speed is compared against the reference motor speed $\omega_{i}$, and the motor failure index is defined as:
@@
When the condition $M_{i}\leq\gamma_{M}$ is satisfied, where $0<\gamma_{M}\leq 1$ denotes the motor failure threshold, the control mode is switched to FTC mode.
@@
When $P_{i}>\gamma_{P}$ where $0<\gamma_{P}\leq 1$ serves as the trigger threshold, the $i$th propeller is classified as failed and isolated. The failure information of the $i$th propeller is subsequently transmitted to trigger controller reconfiguration and trajectory planning.
@@
Beyond complete rotor failures, the FDD framework can also detect the partial thrust reductions (e.g., due to wear, overheating, or intermittent performance) by combining a motor-side cue and a propeller-side cue. Intermittent degradations are detectable if they persist within a short observation window or appear with sufficient frequency. Certain mechanical anomalies such as shaft misalignment typically manifest as speed–thrust inconsistencies and are therefore also within the detectable scope of the proposed FDD. Extremely subtle or very short-lived anomalies may remain undetected, suggesting future extensions that incorporate vibration or thermal cues.
@@ V-C Rotor-Failure-Aware Trajectory Planning
For a polynomial trajectory with $M$ segments in $m$ dimensions and a degree $N_{0}=2s-1$ (where $s=3$ is the integrator chain order), a linear-complexity mapping is defined as:
@@
where $\bm{c}=(\bm{c}_{1}^{T},\ldots,\bm{c}_{M}^{T})^{T}$ represents the polynomial coefficient matrix [1]. The mapping $M(\bm{q}^{p},\bm{T}^{p})$ smoothly relates the intermediate waypoints $\bm{q}^{p}=(\bm{q}_{1},\ldots,\bm{q}_{M-1})^{T}$ and time allocation $\bm{T}^{p}=(T_{1},\ldots,T_{M})^{T}$ to $\bm{c}$. Using $M(\bm{q}^{p},\bm{T}^{p})$, the optimization variables are transformed into $\{\bm{q}^{p},\bm{T}^{p}\}$. The overall trajectory $\bm{p}(t)$ and its individual segment $\bm{p}_{g}(t)$ are expressed as:
$\bm{p}(t)=\bm{p}_{g}(t-t_{g-1}),\quad\forall t\in[t_{g-1},t_{g}],$ (20)
$\bm{p}_{g}(t)=\bm{c}_{g}^{T}\bm{\beta}(t),\quad\forall t\in[0,T_{g}],$ (21)
where $\bm{\beta}(t)=[1,t,\ldots,t^{N_{0}}]^{T}$ denotes the natural polynomial basis. The trajectory generation for the quadrotor is formulated as an unconstrained nonlinear optimization problem:
@@
In reference to [1], we impose limits on linear velocity, acceleration, and jerk to ensure dynamic feasibility during trajectory generation. The corresponding cost function is defined as:
@@
$J_{d,v}=\sum_{i=1}^{k}\max\left\{{\dot{p}^{2}(t_{i})}-v_{{max}}^{2},0\right\},$ (26)
$J_{d,a}=\sum_{i=1}^{k}\max\left\{{\ddot{p}^{2}(t_{i})}-a_{{max}}^{2},0\right\},$ (27)
@@
$J_{d,j}=\sum_{i=1}^{k}\max\left\{{\dddot{p}^{2}(t_{i})}-j_{{max}}^{2},0\right\},$ (29)
@@
$a_{{max},f}=\gamma_{a,f}\left|\frac{\sqrt{T_{f,{max}}^{2}-(m_{0}g_{0})^{2}}-f_{{max}}}{m_{0}}\right|,$ (33)
where $0<\gamma_{a,f}\leq 1$ is a safety factor to enhance flight reliability. To ensure safe operation under rotor failure conditions, dynamic feasibility constraints are explicitly considered. The bounds on velocity and jerk are defined in Eqs. (26) and (29), respectively. Their threshold values are selected empirically based on the characteristics of the quadrotor platform and extensive experimental observations, following a similar practice to that in [33, 34]. By contrast, the acceleration constraint updated in Eq. (33) serves as the primary safety limitation in post-failure scenarios, as it directly reflects the reduced actuation capability after rotor degradation. Although we do not explicitly model efficiency losses or battery degradation, our trajectory optimization already includes total time minimization, which serves as a proxy for energy consumption. This reflects the fact that shorter flight times typically correspond to reduced power usage.
@@
$(\bm{x}^{p}-\bm{s}^{p})^{T}\bm{v}^{p}=0,\quad\bm{x}^{p}\in\mathbb{R}^{3},$ (34)
where $\bm{s}^{p}\in\mathbb{R}^{3}$ denotes a point on the plane, and $\bm{v}^{p}\in\mathbb{R}^{3}$ is the normal vector pointing toward the free space. To quantify the relationship between a given point $\bm{p}^{p}\in\mathbb{R}^{3}$ and the obstacle, the distance $d_{o}$ to the plane is computed as:
$d_{o}=(\bm{p}^{p}-\mathbf{s}^{p})^{T}\mathbf{v}^{p}.$ (35)
@@
$J_{c}=\sum_{i=0}^{\kappa}\max\left\{\left(C_{o}-d_{o}(p(t_{i}))\right),0\right\}^{3},$ (36)
where $C_{o}$ denotes obstacle clearance and $d_{o}$ depends on the position.
@@
Figure 6: Simulation visualization [2] with the proposed method in Test #1, where the propeller-ejection failure occurs during quadrotor hovering. The green and red paths indicate the reference lemniscate trajectory and the real flight trajectory, respectively.